### v1

# SiNGER: A Clearer Voice DistillsVision Transformers Further

Abstract
Vision Transformers are widely adopted as the backbone of vision foundation models, but they are known to produce high-norm artifacts that degrade representation quality. When knowledge distillation transfers these features to students, high-norm artifacts dominate the objective, so students overfit to artifacts and underweight informative signals, diminishing the gains from larger models. Prior work attempted to remove artifacts but encountered an inherent trade-off between artifact suppression and preserving informative signals from teachers. To address this, we introduce Singular Nullspace-Guided Energy Reallocation (SiNGER), a novel distillation framework that suppresses artifacts while preserving informative signals. The key idea is principled teacher feature refinement: during refinement, we leverage the nullspace-guided perturbation to preserve information while suppressing artifacts. Then, the refined teacher’s features are distilled to a student. We implement this perturbation efficiently with a LoRA-based adapter that requires minimal structural modification. Extensive experiments show that SiNGER consistently improves student models, achieving state-of-the-art performance in multiple downstream tasks and producing clearer and more interpretable representations.
**footnotetext: These authors contributed equally to this work.
(a) Overview of SiNGER distillation.
(b) Performance gains using SiNGER.
Figure 1: SiNGER suppresses artifacts and enhances transfer. (a) Feature visualizations highlight clearer and more interpretable representations. (b) Radar chart shows consistent multi-task gains.
1 Introduction
Transformers have become the de facto standard architecture in both research and industry due to their scalability and effectiveness (dinov2; clip). Their token-based self-attention mechanism is broadly applicable with minimal inductive bias (transformer-inductive-bias), and has enabled significant advances in computer vision and machine learning (openvla; unic). Vision Transformers (ViTs, vit) extend this paradigm to visual data and form the backbone of Vision Foundation Models (VFMs). Compared to convolutional networks, ViTs rely less on spatial inductive biases and instead exploit scale to achieve high performance. ViTs are also highly scalable since supervised or self-supervised training produces increasingly generalizable representations (dinov2; deit3).
FitNet (kd-feature-01)
ViTKD (kd-vitkd)
SiNGER (ours)
Input
ViT-Large
ViT-Tiny
FitNet
ViTKD
SiNGER
Figure 2: Qualitative analysis. Row 1: KD method comparison. Left: distilled feature map colored by patch norm, Right: patch-wise cosine similarity to the teacher. Row 2: Input image, two pretrained ViTs, and three ViT-L $\to$} ViT-T distilled variants. Each panel shows similarity from the $\times$-marked patch. SiNGER most closely preserves teacher semantics, showing the most coherent teacher-consistent similarity patterns.
However, the quadratic complexity of self-attention severely limits the practicality of scaling ViTs. This tension between accuracy and efficiency motivates the study of compression. Pruning (vit-pruning-01) and quantization (vit-quant-01) have been explored, but pruning often fails to deliver practical speedup due to structural rigidity (pruning-not-flexible), and quantization can induce numerical instability (quant-not-general-01; quant-not-general-02). Knowledge distillation (KD, kd-hinton) has emerged as the most reliable solution for transferring knowledge from large ViTs to smaller students. KD methods span diverse targets and frameworks (kd-response-01; kd-feature-01; kd-multi-teacher-01), consistently yielding structurally and numerically stable compact models.
Nevertheless, KD for ViTs suffers from subtle but critical limitations in their representation space. vit-needs-registers revealed that ViT token representations contain high-norm artifacts. gandikota2023sinder argue that these artifacts are singular defects induced by power-iteration-like accumulation across residual blocks, whereby tokens align with the leading left singular vector of the pre-trained weights. These artifacts interact poorly with the standard feature mean squared error objective in KD: when the teacher and student are matched, gradients concentrate on the few high-norm tokens, producing an outlier-driven optimization bias that obscures informative signals in the inlier structure. Therefore, suppressing outlier norms in teacher features is essential for KD in ViTs as the scale grows. Prior work mitigated this issue via random masking of teacher features (kd-vitkd); however, this inevitably removes informative signals. Therefore, a key challenge is to mitigate these artifacts without losing valuable information, a fundamental trade-off that requires a principled approach.
To resolve this trade-off, we introduce a nullspace-guided suppression: we modify only the nullspace component in the teacher features, mathematically, the subspace orthogonal to the downstream space. This yields student-optimal supervision by suppressing artifacts without sacrificing informative signals. Based on this insight, we propose Singular Nullspace-Guided Energy Reallocation (SiNGER), a framework that addresses this trade-off in ViTs distillation, illustrated in Figure 1(a). To minimize the modification to the teacher’s signal, we attach a lightweight LoRA-based adapter (hu2021lora) to the KD architecture, which refines the teacher features. The adapter produces a minimal perturbation guided toward the left-nullspace of the next block, suppressing high-norm outliers while leaving the next block output unchanged. Our method achieves superior performance compared to baselines across multiple downstream tasks (Figure 1(b)). It also produces more structured and interpretable feature maps (Figure 2). Our contributions are summarized as follows:
  • We propose a novel distillation framework (SiNGER) that refines teacher signals via the LoRA-based adapter with nullspace initialization to guide effective perturbations.
  • We analyze a fundamental limitation of naïve ViT distillation, showing degraded transfer on downstream benchmarks along with qualitative evidence.
  • We provide extensive ablation studies to analyze the contribution of each component in SiNGER and validate the robustness of our framework.
  • We demonstrate through extensive experiments that our method exceeds baseline performance across tasks and produces more interpretable feature maps.
2 Related Works
Vision Transformers. ViTs (vit) underpin many VFMs and have become the representative architecture for large-scale visual learning. Unlike convolutional networks (resnet; mobilenet), ViTs rely on self-attention with fully connected layers. Since ViTs form the architectural core of most VFMs, studying them provides representative insights that generalize broadly. Similarly, parameter-efficient tuning methods such as LoRA (hu2021lora) highlight how minimal perturbations can effectively adapt VFMs, a perspective that motivates our artifact-suppressing perturbations. However, the quadratic complexity of the ViT architecture limits the practicality of large models despite their advanced representation.
Knowledge Distillation. KD compresses models by training a smaller student to mimic a larger teacher (kd-hinton). Among various approaches, FitNet-style methods (kd-feature-01) that align intermediate features are especially influential, as they encourage the student to learn useful representations beyond logits (kd-response-01). Later extensions incorporated relational structures (kd-relation-01) or multi-teacher settings (kd-multi-teacher-01), while adaptations for ViTs (deit3) aimed to respect their architectural characteristics. Despite these advances, distillation applied to VFMs often inherits undesirable properties from teachers, revealing the need for methods that improve not only compression but also the quality of transferred representations.
Artifacts in Transformers. Artifacts are a recurring issue in transformer models, degrading the representation quality. vit-needs-registers demonstrated that ViTs produce high-norm artifacts, particularly in background regions, harming interpretability and dense prediction. Practical suppression strategies include register tokens (vit-needs-registers), and recent work argues these artifacts arise from power-method-like accumulation across residual layers, aligning tokens with the leading left singular vector  (gandikota2023sinder). In the knowledge distillation domain, ViTKD (kd-vitkd) randomly masks teacher features to reduce the mimicking of high-norm artifacts. However, such indiscriminate masking also removes informative inlier signals, motivating artifact-aware KD that suppresses high-norm artifacts while preserving inlier structure.
3 Method
3.1 Problem Formulation
3.1.1 High-Norm Outliers in Vision Transformers
In large ViTs, a non-negligible fraction of patch features in